In Millikan's experiment, the charge on each drop of oil was measured in coulombs. Imagine the same experiment, but with charges measured in a fictitious unit called a zeet (Z).

Drop Charge (Z)
A 1.44×10−14
B 2.88×10−14
C 3.24×10−14
D 3.60×10−14
E 5.04×10−14

What is the charge on an electron in zeets?
Answer #1

We can do this by subtracting each of the values from the previous value and recording the differences:
(2.88*10^−14 ) - (1.44*10^−14) = 1.44*10^-14
(3.24*10^-14) - (2.88*10^−14) = 3.6*10^-15
(3.60*10^-14) - (3.24*10^-14) = 3.6*10^-15
(5.04*10^-14) - (3.60*10^-14) = 1.44*10^-14
greatest common factor among the 5 measured charges and their differences is = 3.6*10^-15.

So we can hypothesize that 3.6*10^-15 Zeets is the charge of an electron.

charge on an electron in zeets = 3.6*10^-15 Zeets